The following table shows the product price of a perfectly competitive firm and the output produced by the firm. From the table, it can be said that the marginal revenue product of the third machine is _____.
Table 11.1

Machines

Output

Product Price ($)

0

0

3

1

7

3

2

12

3

3

15

3

4

17

3

5

18

3

a.

​$3

b.

​$9

c.

​$45

d.

​3 units

e.

​15 units
Answer #1

Marginal revenue product = marginal product*price of the product

The marginal output of the third output = 15-12 = 3

MRP = 3*3 = 9

option(B)